OpenStudy (anonymous):

how do you solve 13-5x/8 is greater or equal to -4? please show work

OpenStudy (shane_b):

Is that \[13-\frac{5x}{8}=-4\]or is it:\[\frac{13-5x}{8}=-4\]

OpenStudy (anonymous):

second one

OpenStudy (shane_b):

I'm lagging so bad I can't even type...going to try restarting.

OpenStudy (anonymous):

alright

OpenStudy (shane_b):

Ok, let's see how this goes. What do think the first step should be?

OpenStudy (anonymous):

multiply both sides by 8?

OpenStudy (shane_b):

Yes, that leaves you with:\[13-5x>=-32\]Now we'd like to get x by itself so subtract 13 from both sides as well to get:\[-5x>=-45\]

OpenStudy (shane_b):

What do you think is next?

OpenStudy (anonymous):

divide

OpenStudy (shane_b):

Yes, divide both sides by 5 to get:\[-x >=-9\] The final step is to get rid of the negative signs. To do that, you multiply both sides by -1 and flip the sign. That gives you the final answer of:\[x <=9\]

OpenStudy (anonymous):

Thanks

OpenStudy (shane_b):

np